# DefVINS: Visual–Inertial Odometry for Deformable Scenes

### v1

@@ V-NR VI-R Full V-NR VI-R Full V-NR VI-R Full
Based on the previously defined residuals, Fig. 1 illustrates the explicit block structure of the observability matrix $\mathcal{O}$. For clarity and readability, the analytical expressions of the nonzero Jacobian blocks are reported in the text rather than embedded in the figure. Geometric reprojection residuals contribute standard visual–inertial Jacobian blocks, namely $\mathbf{H}_{\mathtt{R}_{t}}^{\mathrm{vis}}=-\mathbf{J}_{\pi}\,\mathbf{p}_{\mathrm{cam}}^{\wedge}$, $\mathbf{H}_{\mathbf{p}_{t}}^{\mathrm{vis}}=-\mathbf{J}_{\pi}\mathtt{R}_{t}^{\top}$, and $\mathbf{H}_{\mathrm{NR}}^{\mathrm{vis}}=\mathbf{J}_{\pi}\mathtt{R}_{t}^{\top}$ for the deformation nodes. Non-rigid regularization terms introduce additional couplings among deformation variables: the elastic prior yields $\mathbf{H}_{\mathrm{NR}}^{\mathrm{elas}}\propto(\mathbf{x}_{i}^{t}-\mathbf{x}_{j}^{t})/d_{ij}^{t}$, the viscous term results in block structures of the form $\mathbf{I},-\mathbf{I},-\mathbf{I},\mathbf{I}]$, and the photometric term contributes $\mathbf{H}_{\mathrm{NR}}^{\mathrm{photo}}=\nabla I^{t}(\mathbf{u}_{i}^{t})\,\mathbf{J}_{\pi}\mathtt{R}_{t}^{\top}$, with analogous derivatives with respect to the pose variables. Together, these expressions define all nonzero blocks of $\mathcal{O}$.
A symbolic inspection of the resulting Jacobian structure shows that the formulation preserves the classical gauge freedoms of visual–inertial odometry while introducing additional degrees of freedom associated with the non-rigid deformation field. When the non-rigid variables $\boldsymbol{\xi}_{\mathrm{NR}}$ are ignored and metric depth is assumed, the rigid subsystem remains observable only up to a global $SE(3)$ transformation: absolute position and yaw are unobservable, whereas gravity magnitude and direction become observable under sufficiently rich accelerations and rotations [32]. In particular, the IMU velocity and position residuals $\mathbf{r}_{\Delta\mathbf{v}}^{t}$ and $\mathbf{r}_{\Delta\mathbf{p}}^{t}$ couple accelerometer bias and gravity, becoming rank-deficient under near-constant-velocity motion, while insufficient rotational excitation leaves yaw and gyroscope bias coupled in $\mathbf{r}_{\Delta\mathtt{R}}^{t}$.
The non-rigid substate $\boldsymbol{\xi}_{\mathrm{NR}}$ is constrained by the combined action of geometric reprojection, photometric consistency, and visco–elastic regularization. Visual and photometric terms anchor node motion to image evidence, whereas elastic and viscous priors regularize spatial and temporal variations, eliminating low-energy deformation modes that would otherwise be consistent with visual measurements alone. As a result, the rank of $\mathcal{O}$ increases in the non-rigid subspace.
@@
In practice, the rank of $\mathcal{O}$ depends on both motion and deformation patterns. Trajectories with limited parallax or negligible rotation lead to weak observability of inertial quantities and non-rigid modes, whereas well-excited motions combined with visco–elastic regularization yield a well-conditioned system in which the rigid state is observable up to the expected gauge freedoms and the deformation field is significantly constrained.
V Experiments
All experiments were conducted on a desktop PC with an Intel Core i7-11700K (3.6 GHz, 64 GB RAM). The proposed method is implemented in C++ and optimized using Ceres [^1] http://ceres-solver.org/. Unless otherwise stated, all methods run in a single thread and wall-clock times are reported.
Performance is evaluated using standard trajectory-based metrics. We report Absolute Trajectory Error (ATE) and Absolute Rotation Error (ARE) to assess global accuracy, Relative Pose Error (RPE) in translation and rotation to measure drift, and the number of successfully tracked frames as an indicator of robustness. Comparisons are performed against ORB-SLAM3 [12] as a rigid SLAM baseline and NR-SLAM [19] as a non-rigid alternative.
We first evaluate the method on the Drunkard’s Dataset [33], which provides 19 synthetic sequences ($\times320$) with full 3D ground truth and four increasing deformation levels per scene, enabling controlled analysis. We then validate the approach on seven real RGB–D sequences ($8×\times4$80) recorded in an industrial setup following [34], using synchronized cameras, IMU, and ground-truth motion. The sequences (R0–R6) capture progressively stronger non-rigid deformations of a textured mandala cloth.
V-A Synthetic experiments
@@
To evaluate the robustness of the proposed initialization strategies under controlled deformation conditions, synthetic experiments are conducted on the Drunkard’s Dataset. This dataset consists of synthetic RGB-D sequences exhibiting progressively increasing levels of non-rigid deformation, ranging from near-rigid motion to large-amplitude surface dynamics. Since inertial measurements are not provided, IMU data are synthetically generated by differentiating the ground-truth trajectories represented as B-splines, ensuring temporally smooth and physically consistent signals. Table II reports quantitative results on the Drunkard’s Dataset, which comprises 19 synthetic scenes evaluated at four increasing levels of deformation difficulty. Results are averaged across scenes at each deformation level to provide a compact yet representative comparison. In low-deformation scenarios (L0), all methods achieve very low errors, with differences in ATE remaining below approximately 35%. In this regime, the rigid visual–inertial baseline ORB-SLAM3 performs competitively with the proposed approaches, indicating that explicit non-rigid modeling is not strictly necessary when deformations are minimal. Nevertheless, the incorporation of inertial measurements already provides measurable benefits: compared to the visual-only DefVINS configuration (V-NR), the rigid visual–inertial variant (VI-R) reduces ATE by roughly 20%. As deformation increases to moderate levels (L1), performance differences become more pronounced. While ORB-SLAM3 exhibits a substantial increase in ATE relative to L0, both NR-SLAM and DefVINS maintain significantly lower errors. In particular, the full DefVINS formulation achieves an ATE reduction of approximately 30% with respect to ORB-SLAM3 and nearly 45% compared to the visual-only baseline.
@@
Improvements in RPE remain more moderate (around 20–30%), suggesting that inertial constraints primarily contribute to stabilizing short-term motion estimates. In harder deformation regimes (L2), rigid motion assumptions further degrade estimation accuracy. Relative to ORB-SLAM3, the full DefVINS formulation reduces ATE by approximately 35–40% and RPE by about 25%. Moreover, when compared to the rigid DefVINS variant (VI-R), the inclusion of explicit non-rigid regularization yields an additional ATE reduction of roughly 30%, highlighting the importance of modeling deformation dynamics beyond inertial stabilization alone. In the extreme deformation case (L3), the limitations of rigid models become most evident. The full DefVINS formulation achieves an ATE reduction of approximately 40–45% with respect to ORB-SLAM3 and nearly 50% compared to the visual-only configuration, while also reducing RPE by about 40%. Although all methods experience a reduction in the number of successfully tracked frames as deformation severity increases, DefVINS consistently maintains longer trajectories, indicating improved robustness to severe non-rigid motion.
@@
TABLE III: Comparison of visual–inertial odometry methods on our real deformable sequences. We report ATE RMSE, translational RPE, and number of successfully tracked frames. Best result per metric in bold. Last row represents the mean for each column.
@@ V-NR VI-R Full V-NR VI-R Full V-NR VI-R Full
Figure 3: Qualitative comparison of DefVINS operating modes and baselines under R4 sequence. a) Rigid VI SLAM (ORB-SLAM3): a representative state-of-the-art rigid visual–inertial system, which may suffer from tracking loss and relocalization in deformable scenes. b) Non-rigid SLAM (NR-SLAM): a representative state-of-the-art for non-rigid environments, which also suffer from tracking loss and relocalization due to its focus on medical datasets. c) DefVINS-V (visual-only, non-rigid): absence of inertial constraints leads to accumulated drift, particularly during turning motions. b) DefVINS-VI (rigid): introducing inertial sensing stabilizes rotational estimates, but assuming scene rigidity limits performance under deformation. e) DefVINS (full): by jointly enabling inertial sensing and explicit non-rigid modeling, the proposed system achieves the highest global consistency and lowest trajectory error. Dashed line represent the ground-truth while colored line represent the camera trajectory.
V-B Observability analysis
@@
To complement the quantitative evaluation, we present an illustrative analysis of numerical observability under controlled synthetic conditions. A set of short motion segments with rich 6-DoF excitation is selected from the synthetic dataset. Specifically, we extract five non-overlapping segments of ten consecutive keyframes each, characterized by non-collinear translations and rotations about multiple axes. For each segment, residual Jacobians are generated and stacked over $k$ consecutive keyframe pairs to form the observability matrix $\mathcal{O}$. The reported results correspond to the average conditioning score across the selected segments. Numerical observability is quantified using the conditioning score $\rho_{k}={\sigma_{\min}}/{\sigma_{\max}}$, where $\sigma_{\min}$ and $\sigma_{\max}$ denote the smallest and largest singular values of $\mathcal{O}$, respectively. Fig. 2 shows $\log_{10}(\rho_{k})$ as a function of the number of stacked keyframe pairs for the visual-only, rigid visual–inertial, and full formulations. The numerical conditioning evolves very differently for the aforementioned configurations. Inertial constraints lift near-null modes associated with gravity and biases, and non-rigid regularization prevents deformation variables from absorbing rigid-body drift. As a result, the full formulation becomes well conditioned with only a few frames, explaining the improved robustness observed in the synthetic experiments.
V-C Realistic experiments
@@
The synthetic evaluation is complemented with a set of realistic experiments aimed at assessing the behavior of the proposed non-rigid regularization under natural image noise, real sensor motion, and uncontrolled surface deformations. To this end, a dedicated dataset composed of seven RGB-D sequences is employed, each captured on a textured deformable surface and exhibiting a different degree of non-rigid motion. The sequences are denoted as R0–R6, where R0 corresponds to near-rigid motion and R6 represents the most severe deformation. This gradual increase in deformation enables a systematic evaluation of the stability of the proposed visco–elastic prior under progressively more challenging conditions. Table III summarizes the quantitative performance of all evaluated methods on the real deformable sequences R0–R6. Several consistent trends can be observed across deformation regimes.
@@
For the nearly rigid sequences (R0–R1), the rigid visual–inertial baseline ORB-SLAM3 achieves the lowest ATE and RPE values, outperforming the deformable models by approximately 20–30% in ATE. This confirms that classical rigid visual–inertial approaches remain highly effective when surface deformation is negligible. In this regime, all DefVINS variants are able to track almost the full sequence length, with only marginal accuracy degradation. As the degree of deformation increases to moderate levels (R2–R3), rigid assumptions progressively break down. Although ORB-SLAM3 and NR-SLAM still maintain tracking, both exhibit a marked increase in ATE. In contrast, incorporating inertial measurements within the DefVINS framework (VI-R) reduces ATE by approximately 25–30% with respect to the visual-only configuration (V-NR), while also lowering RPE by around 15–20%. This highlights the stabilizing effect of inertial constraints on rotational estimation and short-term motion consistency. NR-SLAM consistently yields intermediate performance, improving upon rigid ORB-SLAM3 while remaining less robust than deformable formulations. For medium-to-high deformation scenarios (R4–R6), the limitations of rigid motion models become evident. Relative to ORB-SLAM3, the full DefVINS formulation reduces ATE by approximately 75% on R4, around 80% on R5, and close to 80% on R6. At the same time, DefVINS (Full) consistently tracks a substantially larger fraction of frames (around 85–95%) compared to ORB-SLAM3, whose tracking coverage drops below 50% in R4 and below 20% in R6. While inertial sensing alone partially mitigates tracking degradation, only the joint combination of inertial constraints and explicit non-rigid regularization preserves both accuracy and robustness under strong deformation.
@@
To illustrate the estimated surface behavior, Fig. 4 shows the deformation graph for a representative sequence. The graph reveals spatially varying deformation, with larger displacements in the lower-left region and comparatively stable behavior elsewhere, demonstrating the ability of the proposed visco–elastic regularization to capture localized deformations while maintaining global surface coherence.
VI Conclusions
This paper introduced DefVINS, an observability-gated visual–inertial odometry framework for deformable scenes that decouples a rigid, IMU-anchored state from progressively activated non-rigid deformation parameters. Extensive evaluations on both synthetic and real datasets demonstrate that the proposed formulation provides accurate and stable state estimation across a wide range of deformation regimes. By combining IMU anchoring with conditioning-aware activation of deformation degrees of freedom, DefVINS avoids early over-fitting and catastrophic drift, constituting a suitable and reliable solution for visual–inertial odometry in the presence of non-rigid scene dynamics.

### v2

@@ V-NR VI-R Full V-NR VI-R Full V-NR VI-R Full
Based on the previously defined residuals, Fig. 1 illustrates the explicit block structure of the observability matrix $\mathcal{O}$. For clarity and readability, the analytical expressions of the nonzero Jacobian blocks are reported in the text rather than embedded in the figure. Geometric reprojection residuals contribute with standard visual–inertial Jacobian blocks, namely $\mathbf{H}_{\mathtt{R}_{\tau}}^{\mathrm{vis}}=-\mathbf{J}_{\pi}\,\mathbf{t}_{\mathrm{cam}}^{\wedge}$, $\mathbf{H}_{\mathbf{t}_{\tau}}^{\mathrm{vis}}=-\mathbf{J}_{\pi}\mathtt{R}_{t}^{\top}$, and $\mathbf{H}_{\mathrm{nr}}^{\mathrm{vis}}=\mathbf{J}_{\pi}\mathtt{R}_{\tau}^{\top}$ for the deformation nodes. Non-rigid regularization terms introduce additional couplings among deformation variables: the elastic prior yields $\mathbf{H}_{\mathrm{nr}}^{\mathrm{elas}}\propto(\mathbf{x}_{i}^{\tau}-\mathbf{x}_{j}^{\tau})/d_{ij}^{\tau}$, the viscous term results in block structures of the form $\mathbf{I},-\mathbf{I},-\mathbf{I},\mathbf{I}]$, and the photometric term contributes $\mathbf{H}_{\mathrm{nr}}^{\mathrm{photo}}=\nabla I^{\tau}(\mathbf{u}_{i}^{\tau})\,\mathbf{J}_{\pi}\mathtt{R}_{\tau}^{\top}$, with analogous derivatives with respect to the pose variables. Together, these expressions define all nonzero blocks of $\mathcal{O}$.
A symbolic inspection of the resulting Jacobian structure shows that the formulation preserves the classical gauge freedoms of visual–inertial odometry while introducing additional degrees of freedom associated with the non-rigid deformation field. When the non-rigid variables $\{\mathbf{x}_{i}^{\tau}\}_{i\in\mathcal{D},\tau\in\mathcal{W}}$ are ignored and metric depth is assumed, the rigid subsystem remains observable only up to a global $SE(3)$ transformation: absolute position and yaw are unobservable, whereas gravity magnitude and direction become observable under sufficiently rich accelerations and rotations [31]. In particular, the IMU velocity and position residuals $\mathbf{r}_{\Delta\mathbf{v}}^{\tau}$ and $\mathbf{r}_{\Delta\mathbf{t}}^{\tau}$ couple accelerometer bias and gravity, becoming rank-deficient under near-constant-velocity motion, while insufficient rotational excitation leaves yaw and gyroscope bias coupled in $\mathbf{r}_{\Delta\mathtt{R}}^{t}$.
The non-rigid substate is constrained by the combined action of geometric reprojection, photometric consistency, and visco–elastic regularization. Visual and photometric terms anchor node motion to image evidence, whereas elastic and viscous priors regularize spatial and temporal variations, eliminating low-energy deformation modes that would otherwise be consistent with visual measurements alone. As a result, the rank of $\mathcal{O}$ increases in the non-rigid subspace.
@@
In practice, the rank of $\mathcal{O}$ depends on both motion and deformation patterns. Trajectories with limited parallax or negligible rotation lead to weak observability of inertial quantities and non-rigid modes, whereas well-excited motions combined with visco–elastic regularization yield a well-conditioned system in which the rigid state is observable up to the expected gauge freedoms and the deformation field is sufficiently constrained.
IV Experiments
All experiments were conducted on a desktop PC with an Intel Core i7-11700K (3.6 GHz, 64 GB RAM). The proposed method is implemented in C++ and uses Ceres [^1] http://ceres-solver.org/ for optimization purposes. Unless otherwise stated, all methods run in a single thread and wall-clock times are reported. Performance is evaluated using standard trajectory-based metrics. We report Absolute Trajectory Error (ATE), Relative Position Error (RPE), and the number of successfully tracked frames as an indicator of robustness. Comparisons are performed against ORB-SLAM3 [10] as a rigid visual-inertial SLAM baseline and NR-SLAM [23] as a non-rigid visual alternative. We first evaluate the method on the Drunkard’s Dataset [32], which provides 19 synthetic sequences ($\times320$) with full 3D ground truth and four increasing deformation levels per scene, enabling controlled analysis. We then further validate DefVINS on our VIMandala dataset, comprising seven $8×\times4$80 real sequences for which synchronized RGB, depth IMU, and ground-truth camera motion are recorded. The sequences , that we denote as R0–R6, capture progressively stronger non-rigid deformations of a textured mandala cloth.
IV-A Synthetic experiments in the Drunkard’s dataset
To evaluate the robustness of DefVINS under controlled conditions, we evaluate it on the Drunkard’s Dataset. This dataset consists of synthetic RGB-D sequences exhibiting progressively increasing levels of non-rigid deformation, ranging from near-rigid motion to large-amplitude surface dynamics. As inertial measurements are not provided, IMU data are synthetically generated by differentiating ground-truth trajectories represented as B-splines, ensuring temporally smooth and physically consistent signals.
Table I reports quantitative results on Drunkard’s, which comprises 19 synthetic scenes evaluated at four increasing levels of deformation difficulty. Results are averaged across scenes at each deformation level to provide compact yet representative metrics. For our DefVINS, we report results for the full pipeline, but also for the ablated versions V-NR (no inertial measurements) and VI-R (rigid scene model).
@@ IV-A Synthetic experiments in the Drunkard’s dataset
In low-deformation scenarios (L0), all methods achieve similarly low errors. In this regime, the rigid visual–inertial baseline ORB-SLAM3 performs competitively against non-rigid approaches, indicating that explicit non-rigid modeling is not strictly necessary when deformations are minimal. Nevertheless, the incorporation of inertial measurements already provides measurable benefits: compared to the visual-only DefVINS configuration (V-NR), the rigid visual–inertial variant (VI-R) reduces ATE by roughly 20%. As deformation increases to moderate levels (L1), performance differences become more pronounced. While ORB-SLAM3 exhibits a substantial increase in ATE relative to L0, both NR-SLAM and DefVINS maintain significantly lower errors. In particular, our full DefVINS formulation achieves an ATE reduction of approximately 30% with respect to ORB-SLAM3 and nearly 45% compared to the visual-only baseline.
@@
Improvements in RPE remain more moderate (around 20–30%), suggesting that inertial constraints primarily contribute to stabilizing short-term motion estimates. In harder deformation regimes (L2), rigid motion assumptions further degrade estimation accuracy. Relative to ORB-SLAM3, the full DefVINS formulation reduces ATE by approximately 35–40% and RPE by about 25%. Moreover, when compared to the rigid DefVINS variant (VI-R), the inclusion of explicit non-rigid regularization yields an additional ATE reduction of roughly 30%, highlighting the importance of modeling deformation dynamics beyond inertial stabilization alone. In the extreme deformation case (L3), the limitations of rigid models become most evident. The full DefVINS formulation achieves an ATE reduction of approximately 40–45% with respect to ORB-SLAM3 and nearly 50% compared to the visual-only configuration, while also reducing RPE by about 40%. Although all methods experience a reduction in the number of successfully tracked frames as deformation severity increases, DefVINS consistently achieves the longest estimated trajectories, indicating improved robustness to severe non-rigid motion.
@@
TABLE II: Comparison of visual–inertial odometry methods on the VIMandala dataset. We report ATE RMSE, RPE, and number of successfully tracked frames. Best result per metric in bold. Last row shows per-column means.
@@ V-NR VI-R Full V-NR VI-R Full V-NR VI-R Full
Figure 3: Comparison between DefVINS operating modes and baselines in R4 sequence. a) ORB-SLAM3: a representative rigid visual–inertial system. b) NR-SLAM: a representative non-rigid visual SLAM pipeline, with focus on medical applications. c) DefVINS-V (visual-only, non-rigid): the lack of inertial constraints leads to accumulated drift, particularly during turning motions. b) DefVINS-VI (rigid): introducing inertial sensing stabilizes rotational estimates, but assuming scene rigidity limits performance under deformation. e) DefVINS (full): by jointly enabling inertial sensing and explicit non-rigid modeling, DefVINS achieves the global consistency and lowest trajectory error. Dashed line represent the ground-truth while colored line represent the camera trajectory and its estimation errors.
IV-B Observability analysis
@@
To complement the quantitative evaluation, we present an illustrative analysis of numerical conditioning under controlled synthetic conditions. A set of short motion segments with rich 6-DoF excitation is selected from the synthetic dataset. Specifically, we extract five non-overlapping segments of ten consecutive keyframes each, characterized by non-collinear translations and rotations about multiple axes. For each segment, residual Jacobians are generated and stacked over $k$ consecutive keyframe pairs to form the observability matrix $\mathcal{O}$. The reported results correspond to the average conditioning score across the selected segments. Numerical observability is quantified using the conditioning score $\rho_{k}={\sigma_{\min}}/{\sigma_{\max}}$, where $\sigma_{\min}$ and $\sigma_{\max}$ denote the smallest and largest singular values of $\mathcal{O}$, respectively. Fig. 2 shows $\log_{10}(\rho_{k})$ as a function of the number of stacked keyframe pairs for the visual-only, rigid visual–inertial, and full formulations. The numerical conditioning evolves very differently for the aforementioned configurations. Inertial constraints lift near-null modes associated with gravity and biases, and non-rigid regularization prevents deformation variables from absorbing rigid-body drift. As a result, the full formulation becomes well conditioned with only a few frames, explaining the improved robustness observed in the synthetic experiments.
IV-C Real experiments in the VIMandala dataset
@@
The synthetic evaluation is complemented with a set of realistic experiments aimed at assessing the behavior of the proposed non-rigid regularization under natural image noise, real sensor motion, and uncontrolled surface deformations. To this end, we present the VIMandala dataset composed of seven RGB-D sequences, each of them captured on a textured deformable surface and exhibiting a different degree of non-rigid motion. The sequences are denoted as R0–R6, where R0 corresponds to near-rigid motion and R6 represents the most severe deformations. This gradual increase in deformation enables a systematic evaluation of the stability of the proposed visco–elastic prior under progressively more challenging conditions. Table II summarizes the quantitative performance of all evaluated methods on the real deformable sequences R0–R6. Several consistent trends can be observed across deformation regimes.
@@
For the nearly rigid sequences (R0–R1), the rigid visual–inertial baseline ORB-SLAM3 achieves the lowest ATE and RPE values, outperforming the deformable models by approximately 20–30% in ATE. This confirms that classical rigid visual–inertial approaches remain highly effective when surface deformation is negligible, and we attribute these better results to the better engineered implementation of ORB-SLAM3. In this regime, all DefVINS variants are able to track almost the full sequence length, with only marginal accuracy degradation. As the degree of deformation increases to moderate levels (R2–R3), rigid assumptions progressively break down. Although ORB-SLAM3 and NR-SLAM still maintain tracking, both exhibit a marked increase in ATE. In contrast, incorporating inertial measurements within the DefVINS framework (VI-R) reduces ATE by approximately 25–30% with respect to the visual-only configuration (V-NR), while also lowering RPE by around 15–20%. This highlights the stabilizing effect of inertial constraints on rotational estimation and short-term motion consistency. NR-SLAM consistently achieves intermediate performance, outperforming the rigid ORB-SLAM3 but still falling short of our proposed DefVINS. For medium-to-high deformation scenarios (R4–R6), the limitations of rigid scene models become evident. Relative to ORB-SLAM3, the full DefVINS formulation reduces ATE by approximately 75% on R4, around 80% on R5, and close to 80% on R6. At the same time, DefVINS (Full) consistently tracks a substantially larger fraction of frames (around 85–95%) compared to ORB-SLAM3, whose tracking coverage drops below 50% in R4 and below 20% in R6. While inertial sensing alone partially mitigates tracking degradation, only the joint combination of inertial constraints and explicit non-rigid regularization preserves both accuracy and robustness under strong deformation.
@@
To illustrate the estimated surface behavior, Fig. 4 shows the deformation graph for a representative image of our VIMandala dataset. The graph reveals spatially varying deformation, with larger displacements in the lower-left region and comparatively stable behavior elsewhere, demonstrating the ability of the proposed visco–elastic regularization to capture localized deformations while maintaining global surface coherence.
V Conclusions
This paper introduced DefVINS, a visual–inertial odometry pipeline for deformable scenes that decouples a rigid, IMU-anchored state from progressively activated non-rigid deformation parameters. In addition, given the lack of benchmarks for this task, we contribute by synthesizing IMU readings for the synthetic Drunkard’s dataset and curating a new real benchmark imaging a deforming mandala. Our experiments on both synthetic and real data demonstrate that our DefVINS provides accurate and stable state estimation across a wide range of deformation regimes, outperforming relevant rigid visual–inertial and non-rigid visual baselines.